An increase in the market price of mens haircuts, from $20 per haircut to $30 per haircut, initially causes a local barbershop to have its employees work overtime to increase the number of daily haircuts provided from 40 to 45. When the $30 market price remains unchanged for several weeks and all other things remain equal as well, the barbershop hires additional employees and provides 60 haircuts per day. What is the short-run price elasticity of supply?(Your answer should have two decimal places.)
Answer #1

Let us calculate the short-run price elasticity of supply:

The difference between an increased men’s per haircut: $30 - $20 = $10

The difference between an increased number of daily haircuts: 45 – 40 = 5

Short-run price elasticity of supply = Difference of increasing number of daily haircuts / Difference of increase men’s per hair cut.

=5/10

=0.5

Or

=(45-40) / (30-20)

=5/10

=0.5

So, the short-run price elasticity of supply = 0.50.